OpenStudy (anonymous):

what is 1 = 2/3(1) +b ?

OpenStudy (anonymous):

1 = 2/3(1) +b 1 - 2/3 = b b = (3-2)/3 b =1/3

OpenStudy (anonymous):

that is what i thought,,, but the answer is 5/3 :(

OpenStudy (anonymous):

i don't know how to get that

OpenStudy (anonymous):

something is strange, because almost no one write \[1=\frac{3}{4}1+b\] why is the \(1\) there?

OpenStudy (anonymous):

which 1??

OpenStudy (anonymous):

you wrote 1 = 2/3(1) +b that "1" in the parentheses is weird. it does nothing

OpenStudy (anonymous):

it's just part of the problem,,,